OpenStudy (anonymous):

Some scientists estimate the growth of a population using the Malthusian model, named after the Reverend Thomas Malthus, who wrote one of the most influential books on population. His formula uses the natural logarithm, which you will learn about in a future lesson. Another formula that models growth is shown, where No is the initial size of the population, N is the final size of the population, r is the rate of growth or decay per time period, and t is the number of time periods. In this case, r = 100% or 1. Using this formula, how many rabbits would Daisy have after 3 years? N =No(1 + r)t

OpenStudy (anonymous):

i came out with 16384

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@sammixboo

OpenStudy (amistre64):

No is the initial size of the population, N is the final size of the population, r is the rate of growth or decay per time period, and t is the number of time periods. In this case, r = 1. Using this formula, how many rabbits would Daisy have after 3 years? N =No(1 + r)t whats the initial population count, 2 maybe?

OpenStudy (amistre64):

is 3 years broken into smaller periods? or does 1 year measure 1 period?

OpenStudy (anonymous):

(Hints: Remember that your starting number is 2. The exponent is determined by the number of periods of doubling. Use your calculator.)

OpenStudy (anonymous):

sorry forgot to post the first problem

OpenStudy (anonymous):

Daisy purchased a rabbit couple to keep as pets. Perhaps Daisy got more than she bargained for! If the population of rabbits doubles every 3 months, how many rabbits will Daisy have at the end of 3 years? (Hints: Remember that your starting number is 2. The exponent is determined by the number of periods of doubling. Use your calculator.)

OpenStudy (anonymous):

my answer was 8192

OpenStudy (anonymous):

not sure if that helps the 2nd problem alludes me completely lol.

OpenStudy (amistre64):

N = 2(1+1)^3 so long as 1 year is 1 periodic rate. spose they take 3 months to cycle thru 1 period, 4 periods per year is 12 periods in 3 years N = 2(1+1)^12

OpenStudy (anonymous):

ok so N=8192?

OpenStudy (amistre64):

doubles every 3 months gives us n0 = 2 n1 = 2(2) n2 = 2(2)(2) n3 = 2^4 n_i = 2^{i-1}

OpenStudy (amistre64):

2^11 = 2048

OpenStudy (anonymous):

mm you lost me

OpenStudy (amistre64):

was making a geometric progression to make sure of something

OpenStudy (anonymous):

i plugged in N = 2(1+1)^12 and got N=8192

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

No = 2(1+1)^0 N1 = 2(2)^1 N2 = 2(2)^2 ... N12 = 2(2)^12 = 8192 yes

OpenStudy (anonymous):

thank you very much!

OpenStudy (amistre64):

youre welcome